# How In-Context Learning Emerges from Training on Unstructured Data: On the Role of Co-Occurrence, Positional Information, and Noise Structures

### v1

@@ 3.2 In-context learning on dual-pattern tasks
We next examine the case where both the training data and ICL prompts contain two different patterns occurring with equal probability: abcadefd and abcbdefe, where $a\neq b\neq c\neq a$ and $d\neq e\neq f\neq d$. We consider the clean and block-noisy scenarios, defined similarly as in Section 3.1, and set $|V|=N=20$. Table 6 outlines the ICL performance for both scenario types across different model configurations. Unlike the single-pattern scenario, there is an improvement in performance with five layers compared to one layer, particularly with learned positional embeddings.
This phenomenon is related to the notion of induction heads, where at least two layers may be necessary to distinguish the two patterns [Olsson et al., 2022]. This is reflected in Figure 2, which compares the accuracy trajectories of one-layer and five-layer models. While the five-layer setup effectively differentiates the two patterns, the one-layer configuration fails to do so. Meanwhile, in both clean and block-noisy scenarios, learned positional embeddings lead to notably higher accuracies as compared to sinusoidal ones, similar to the single-pattern case.
Scenarios where in-context learning unexpectedly fails
In this section, we consider two scenarios where in-context learning (ICL) unexpectedly fails, irrespective of architectures. In Section 4.1, both the training data and test prompts follow repeating patterns across blocks, but the pattern in the test data differs from that in the training data. In Section 4.2, the training sentences contain known input-output pairs but only at fixed locations. Section 4.3 concludes with a synthetic data experiment that provides support to the theory.
@@ 4.1 Failed scenario 1: Sentences with repeating patterns
In this scenario, the training data comprises sentences in the form of abacdcefe, where $a\neq b$, $c\neq d$, and $e\neq f$. Each sentence is composed of three blocks, and each block consists of three tokens with the same pattern. For the ICL task, we consider predicting $\underline{f}$ from the prompt $\underline{abbcddef}$, where $\underline{a}\neq\underline{b}$, $\underline{c}\neq\underline{d}$, and $\underline{e}\neq\underline{f}$. As each training sentence contains a repeated pattern, we expect a well-trained model to output $\underline{f}$ to maintain the pattern seen in the in-context examples: $\underline{a}\underline{b}\underline{b}$ and $\underline{c}\underline{d}\underline{d}$. However, as depicted in Table 7, all models fail to recognize the repeated patterns and predict the correct token.
We next formalize a generalization of this scenario. Let the vocabulary be $\mathcal{V}=\{1,2,\cdots,|V|\}$, and define $S=\{(a,b)\leavevmode\nobreak\ |\leavevmode\nobreak\ a,b\in\mathcal{V},a\neq b\}$. To ensure training sentences are distinct from the ICL prompts, we first partition $S$ into $S_{1}$ and $S_{2}$, where $\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{1}\}=\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{2}\}=\mathcal{V}$. Here, $c[i]$ denotes the $i$-th element of $c$. Suppose we autoregressively train a sufficiently large position-aware model so that it is possible to achieve the minimum possible theoretical loss. The training sentences take the form $x_{11}x_{12}x_{11}x_{21}x_{22}x_{21}\cdots x_{N1}x_{N2}x_{N1}$, where $x_{i1}\neq x_{i2}$ and $(x_{i1},x_{i2})$ is independently selected from $S_{1}$ for every $i\in[N]$. Theorem 6, whose proof is in to Appendix F, states that ICL fails to hold regardless of the number of in-context examples.
@@
Theorem 6 and Table 7 demonstrate that ICL achieves zero accuracy irrespective of the number of in-context examples ($\ell-1$). This finding provides insight into the generalization capacity of autoregressive models in the context of ICL. Simply put, if the pattern in the in-context examples differs significantly from any pattern in the training data, ICL may not occur.
@@
4.2 Failed scenario 2: Sentences with known pairs but only at fixed locations
We revisit the paired relationship scenario discussed in Section 2. The training data now comprises sentences of the form of $a_{i}pqrsb_{i}$, where $(a_{i},b_{i})$ represents a known pairing and $p,q,r,s$ represent other words. For the ICL task, we consider predicting $b_{i_{3}}$ from the prompt $a_{i_{1}}b_{i_{1}}a_{i_{2}}b_{i_{2}}a_{i_{3}}$, where $i_{1}\neq i_{2}\neq i_{3}\neq i_{1}$. As each training sentence contains an $(a_{i},b_{i})$ pair always at a fixed location, we expect a well-trained model to output $b_{i_{3}}$ to maintain the pattern in the in-context examples: $a_{i_{1}}b_{i_{1}}$ and $a_{i_{2}}b_{i_{2}}$. However, none of the models can identify the repeated patterns and predict the correct token, as shown in Table 7.
We next formalize a generalization of this scenario. Let the vocabulary be $\{(a_{i},b_{i})\}_{i\in[I]}\cup\mathcal{V}$, where $\mathcal{V}=\{1,2,\cdots,|V|\}$ represent other words. As in Section 4.1, we autoregressively train a sufficiently large position-aware model that can achieve the minimum possible theoretical loss. The training sentences take the form $a_{i}v_{1}v_{2}\cdots v_{2k}b_{i}$, where $i$ and $v_{1:2k}$ are independently chosen from $[I]$ and $\mathcal{V}$, respectively, uniformly at random. Theorem 7, whose proof is in Appendix G, states that ICL fails to occur regardless of the number of in-context examples.
@@ 4.2 Failed scenario 2: Sentences with known pairs but only at fixed locations
Consider the generalized scenario in Section 4.2. For any $11\leq\ell\leq k+$ 1, given an in-context prompt of the form $a_{i_{1}}b_{i_{1}}a_{i_{2}}b_{i_{2}}\cdots a_{i_{\ell}}$ with distinct $i_{j}$’s, the model never predicts $b_{i_{\ell}}$.
Theorem 7 highlights the finding that the success of ICL relies heavily on how the patterns appear in the training data. In this scenario, the $(a_{i},b_{i})$ pairs consistently appear at the beginning and end of each training sentence, and we anticipate the model to recognize this relationship for ICL to occur. However, as shown in Theorem 7 and Table 7, this is not the case.
@@ 4.3 Experiment on a synthetic corpus
We conduct an experiment on a synthetic corpus consisting of (country)-(capital) relationships. Each sentence in the corpus is categorized into exactly one of four possible categories: (1) exactly one country-capital pair; (2) exactly two country-capital pairs; (3) exactly one country without any pair; and (4) no country. In sentences with exactly one country-capital pairs, each capital appears in the first position, each country appears in the last position, and every sentence consists of six words. The corpus generation process is as follows:
  1. 1.
  2. 2.
    Generate 130 sentences containing exactly one country-capital pair (13 for each country). Example: Paramaribo stands as capital of Suriname.
  3. 3.
  4. 4.
    Generate 60 sentences without any country, capital city, or IOC code.
  5. 5.
@@
We train a five-layer two-head autoregressive transformer on this corpus, with an embedding dimension of $100$. Similar to Section 2.5, we assess the ICL accuracies using prompts involving countries and their capitals. We discover that the ICL accuracies are zero regardless of the number of in-context examples (one to five), thus supporting the theory.
Related work
@@
Large language models (LLMs), such as transformers, are widely recognized for their outstanding performance in in-context learning (ICL) [Brown et al., 2020]. ICL refers to the capability of LLMs to discern specific tasks and generate predictions based on input-output pairs (known as prompts) without needing any parameter updates. A multitude of studies have been dedicated to exploring this intriguing phenomenon from various theoretical and empirical perspectives. In this section, we provide a brief summary of some of these studies.
@@
Some studies adopted a Bayesian approach to studying ICL. Xie et al. [2021] posited that ICL can be viewed as implicit Bayesian inference. They demonstrated that LLMs can infer a latent document-level concept for next-token prediction during pretraining and a shared latent concept across input-output pairs in an ICL prompt, under the assumption that documents are generated from hidden Markov models (HMMs). Wang et al. [2023] and Zhang et al. [2023] expanded on this idea by exploring more realistic latent variable models beyond HMMs. Wang et al. [2023] argued that large language models function as latent variable models, with latent variables containing task-related information being implicitly inferred. Zhang et al. [2023] showed that without updating the neural network parameters, ICL can be interpreted as Bayesian model averaging parameterized by the attention mechanism. Panwar et al. [2023] provided empirical evidence that transformers behave like Bayesian predictors when performing ICL with linear and non-linear function classes. Dalal and Misra [2024] proposed a Bayesian learning framework to understand ICL through the lens of text generation models represented by multinomial transition probability matrices. Chiang and Yogatama [2024] proposed the pelican soup framework to explain ICL without relying on latent variable models. This framework incorporates concepts such as a common sense knowledge base, natural language classification, and meaning association, enabling the establishment of a loss bound for ICL that depends on the number of in-context examples.
@@
In a related exploration, Li et al. [2023a] showed that softmax regression models learned through gradient descent are similar to transformers. Ren and Liu [2023] related ICL with softmax transformers to contrastive learning, where the inference process of ICL can be viewed as a form of gradient descent. Mahankali et al. [2023] proved that minimizing the pretraining loss is equivalent to a step of gradient descent in single-layer linear transformers. Vladymyrov et al. [2024] established that linear transformers execute a variant of preconditioned gradient descent by maintaining implicit linear models. On the other hand, some studies argued that the ICL ability of transformers cannot be attributed to gradient descent. Fu et al. [2023] showed that ICL for linear regression tasks arises from higher-order optimization techniques like iterative Newton’s method rather than gradient descent. Wibisono and Wang [2023] demonstrated that transformers can perform ICL on unstructured data that lack explicit input-output pairings, with softmax attention playing an important role especially when using a single attention layer. Shen et al. [2023] provided empirical evidence that the equivalence between gradient descent and ICL might not be applicable in real-world scenarios.
@@
Numerous studies focused on the pretraining aspects (e.g., data distribution and task diversity) of ICL. Min et al. [2022] showed that the input-label mapping in the in-context examples does not significantly affect ICL performance. Chan et al. [2022] demonstrated that the ICL capabilities of transformers depend on the training data distributions and model features. Kossen et al. [2024] established that ICL considers in-context label information and is capable of learning entirely new tasks in-context. Li and Qiu [2023] introduced an iterative algorithm designed to enhance ICL performance by selecting a small set of informative examples that effectively characterize the ICL task. Qin et al. [2023] proposed a method based on zero-shot chain-of-thought reasoning for selecting ICL examples, emphasizing the importance of choosing diverse examples that are strongly correlated with the test sample. Han et al. [2023b] studied ICL by identifying a small subset of the pretraining data that support ICL via gradient-based methods. They discovered that this supportive pretraining data typically consist of more uncommon tokens and challenging examples, characterized by a small information gain from long-range context. Peng et al. [2024] proposed a selection method for ICL demonstrations that are both data-dependent and model-dependent. Van et al. [2024] introduced a demonstration selection method that enhances ICL performance by analyzing the influences of training samples using influence functions.
@@
In a similar vein, Wu et al. [2023] demonstrated that pretraining single-layer linear attention models for ICL on linear regression with a Gaussian prior can be effectively accomplished with a minimal number of independent tasks, regardless of task dimension. Raventós et al. [2023] emphasized a task diversity threshold that differentiates the conditions under which transformers can successfully address unseen tasks. Yadlowsky et al. [2023] attributed the impressive ICL capabilities of transformers to the diversity and range of data mixtures in their pretraining, rather than their inductive biases for generalizing to new tasks. Ding et al. [2024] compared the ICL performance of transformers trained with prefixLM (where in-context samples can attend to all tokens) versus causalLM (where in-context samples cannot attend to subsequent tokens), finding that the latter resulted in poorer ICL performance. Chen et al. [2024b] discovered that the ICL capabilities of language models rely on the presence of pairs of phrases with similar structures within the same sentence. Zhao et al. [2024] proposed a calibration scheme that modifies model parameters by adding random noises, resulting in fairer and more confident predictions. Abbas et al. [2024] demonstrated that the ICL predictions from transformer-based models often exhibit low confidence, as indicated by high Shannon entropy. To address this issue, they introduced a straightforward method that linearly calibrates output probabilities, independent of the model’s weights or architecture.
@@
A number of works delved into specific data generating processes to provide insight into the emergence of ICL. Bhattamishra et al. [2023] examined the ICL ability of transformers by focusing on discrete functions. Specifically, they showed that transformers perform well on simpler tasks, struggle with more complex tasks, and can learn more efficiently when provided with examples that uniquely identify a task. Guo et al. [2023] investigated ICL in scenarios where each label is influenced by the input through a potentially complex yet constant representation function, coupled with a unique linear function for each instance. Akyürek et al. [2024] studied ICL of regular languages produced by random finite automata. They compared numerous neural sequence models and demonstrated that transformers significantly outperform RNN-based models because of their ability to develop n-gram heads, which are a generalization of induction heads. Sander et al. [2024] analyzed simple first-order autoregressive processes to gain insight into how transformers perform ICL to predict the next tokens.
@@
Some studies explored how different components of transformers affect their ICL abilities. Ahuja and Lopez-Paz [2023] compared the ICL performance of transformers and MLP-based architectures under distribution shifts. Their findings demonstrate that while both methods perform well in in-distribution ICL, transformers exhibit superior ICL performance when faced with mild distribution shifts. Collins et al. [2024] showed that softmax attention outperforms linear attention in ICL due to its ability to calibrate its attention window to the Lipschitzness of the pretraining tasks. Xing et al. [2024] focused on linear regression tasks to identify transformer components that enable ICL. They found that positional encoding is crucial, along with the use of multiple heads, multiple layers, and larger input dimensions. Cui et al. [2024] proved that multi-head attention outperforms single-head attention in various practical scenarios, including those with noisy labels and correlated features. Chen et al. [2024a] investigated the ICL dynamics of a multi-head softmax attention model applied to multi-task linear regression. They proved the convergence of the gradient flow and observed the emergence of a task allocation phenomenon, where each attention head specializes in a specific task.
@@
Finally, several studies proposed various hypotheses on the emergence of ICL and provided theoretical justifications. Li et al. [2023b] viewed ICL as an algorithm learning problem where a transformer implicitly constructs a hypothesis function at inference time. Han et al. [2023a] argued that the ability of transformers to execute ICL is attributable to their capacity to simulate kernel regression. Singh et al. [2023] explored the interaction between ICL and in-weights learning (IWL) using synthetic data designed to support both processes. They observed that ICL initially emerges, followed by a transient phase where it disappears and gives rise to IWL. Yan et al. [2023] studied ICL from the perspective that token co-occurrences play a crucial role in guiding the learning of surface patterns that facilitates ICL. Abernethy et al. [2024] showed that transformers can execute ICL by dividing a prompt into examples and labels, then employing sparse linear regression to deduce input-output relationships and generate predictions. Lin and Lee [2024] developed a probabilistic model that can simultaneously explain both task learning and task retrieval aspects of ICL. Here, task learning refers to the ability of language models to identify a task from in-context examples, while task retrieval pertains to their ability to locate the relevant task within the pretraining data.
Discussion
In this paper, we investigate how in-context learning (ICL) can emerge from pretraining on unstructured natural language data. We present three main findings, supported by both theory and empirical studies. First, ICL can be achieved by simply modeling co-occurrence using older generations of language models like continuous bag of words (CBOW), when ICL prompts involve pairs that frequently appear together. Second, when ICL prompts involve recognizable patterns that do not always co-occur, positional information and noise structures play crucial roles in enabling ICL. Finally, we highlight the importance of training data structure in ICL by examining two instances where ICL unexpectedly fails.
Limitations and future work
This study has several limitations. Firstly, the experiments are conducted on a relatively small scale. However, they still provide sufficient evidence to support the theoretical findings. Secondly, the focus of this study is on two specific types of in-context learning (ICL) tasks, as described in Section 1. Lastly, real data sets are not utilized due to the lack of alignment with the study objectives. Despite these limitations, we believe that this work offers valuable insights into the emergence of ICL through training on unstructured natural language data, supported by both theoretical and empirical evidence from experiments involving prompting and synthetic data. Further analyses on other ICL tasks and their reliance on model architecture can be fruitful avenues for future work.
Acknowledgments.   This work was supported in part by the Office of Naval Research under grant number N00014-23-1-2590 and the National Science Foundation under Grant No. 2231174 and No. 2310831.
Appendix A Proof of Theorem 1

### v2

@@ 3.2 In-context learning on dual-pattern tasks
We next examine the case where training data and ICL prompts contain two different patterns occurring equally likely: abcadefd and abcbdefe, where $a,b,c$ and $d,e,f$ are distinct. We consider the clean and block-noisy scenarios as in Section 3.1, and set $|V|=N=20$ (details in Appendix J).
Results. Table 6 outlines the ICL performance for both scenario types across different model configurations. Unlike the single-pattern scenario, there is an improvement in performance with five layers compared to one layer, particularly with learned positional embeddings.
This phenomenon is related to the notion of induction heads, where at least two layers may be necessary to distinguish the two patterns [38]. This is reflected in Figure 2, which compares the accuracy trajectories of one-layer and five-layer models. While the five-layer setup effectively differentiates the two patterns, the one-layer configuration fails to do so. Meanwhile, in both clean and block-noisy scenarios, learned positional embeddings lead to notably higher accuracies as compared to sinusoidal ones, similar to the single-pattern case.
Scenarios where in-context learning fails
In this section, we consider two scenarios where in-context learning (ICL) fails, irrespective of architectures. In Section 4.1, we consider a logic reasoning task requiring identification and generalization of a repetition meta-pattern within sequences. In Section 4.2, we explore a word analogy task where relevant word pairs appear in unstructured training sentences but are limited to fixed positions. Section 4.3 concludes with a synthetic data experiment supporting the theory.
@@ 4.1 Failed scenario 1: Sentences with repeating patterns
In this meta-pattern recognition and generalization task, each training sequence follows a repeating pattern based on its starting tokens, and the ICL task sequence requires the model to identify this repetition and extend it to a new, unseen starting pattern. Specifically, our training data comprises sentences in the form of abacdcefe, where $a\neq b$, $c\neq d$, and $e\neq f$. Note that each sentence is structured into three blocks, each consisting of three tokens with the same pattern. For the ICL task, we consider predicting $\underline{f}$ from the prompt $\underline{abbcddef}$, where $\underline{a}\neq\underline{b}$, $\underline{c}\neq\underline{d}$, and $\underline{e}\neq\underline{f}$. Given the repeated pattern within each training sequence, a well-trained model might be expected to output $\underline{f}$ to continue the pattern established in the in-context examples: $\underline{a}\underline{b}\underline{b}$ and $\underline{c}\underline{d}\underline{d}$. However, as seen in Table 7, all models fail to recognize and apply the pattern, resulting in incorrect predictions.
Formalization. We now formalize a generalization of this scenario. Let the vocabulary be $\mathcal{V}=\{1,2,\cdots,|V|\}$, and define $S=\{(a,b)\leavevmode\nobreak\ |\leavevmode\nobreak\ a,b\in\mathcal{V},a\neq b\}$. To ensure training sentences are distinct from the ICL prompts, we first partition $S$ into $S_{1}$ and $S_{2}$, where $\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{1}\}=\{c[1]\leavevmode\nobreak\ |\leavevmode\nobreak\ c\in S_{2}\}=\mathcal{V}$. Here, $c[i]$ denotes the $i$-th element of $c$. Suppose we autoregressively train a sufficiently large position-aware model so that it is possible to achieve the minimum possible theoretical loss. The training sentences take the form $x_{11}x_{12}x_{11}x_{21}x_{22}x_{21}\cdots x_{N1}x_{N2}x_{N1}$, where $x_{i1}\neq x_{i2}$ and $(x_{i1},x_{i2})$ is independently selected from $S_{1}$ for every $i\in[N]$. Theorem 6, whose proof is in Appendix H, states that ICL fails regardless of the number of in-context examples.
@@
Results. Theorem 6 and Table 7 demonstrate that ICL achieves zero accuracy irrespective of the number of in-context examples ($\ell-1$). This insight sheds light on the ICL capacity of autoregressive models. Simply put, if the pattern in the in-context examples differs significantly from any pattern in the training data, ICL may not occur. These results align with the findings of Raventós et al. [42] and Yadlowsky et al. [63] on the importance of data diversity for ICL.
4.2 Failed scenario 2: Sentences with co-occurring word pairs restricted to fixed locations
@@
We revisit the word analogy task in Section 2. The training data now comprises sentences of the form of $a_{i}pqrsb_{i}$, where $(a_{i},b_{i})$ represents a frequently co-occurring word pair and $p,q,r,s$ represent other words. For the ICL task, we consider predicting $b_{i_{3}}$ from the prompt $a_{i_{1}}b_{i_{1}}a_{i_{2}}b_{i_{2}}a_{i_{3}}$, where $i_{1},i_{2},i_{3}$ are distinct. As each training sentence always contains an $(a_{i},b_{i})$ pair at a fixed location, we expect a well-trained model to output $b_{i_{3}}$ to maintain the pattern in in-context examples: $a_{i_{1}}b_{i_{1}}$ and $a_{i_{2}}b_{i_{2}}$. Yet Table 7 shows none of the models can identify the patterns and predict the correct token.
Formalization. We now formalize a generalization of this scenario. Let the vocabulary be $\{(a_{i},b_{i})\}_{i\in[I]}\cup\mathcal{V}$, where $\mathcal{V}=\{1,2,\cdots,|V|\}$ represent other words. As in Section 4.1, we autoregressively train a sufficiently large position-aware model that can achieve the minimum possible theoretical loss. The training sentences take the form $a_{i}v_{1}v_{2}\cdots v_{2k}b_{i}$, where $i$ and $v_{1:2k}$ are independently chosen from $[I]$ and $\mathcal{V}$, respectively, uniformly at random. Theorem 7, whose proof is in Appendix I, states that ICL fails regardless of the number of in-context examples.
@@
Consider the generalized scenario in Section 4.2. For any $11\leq\ell\leq k+$ 1, given an in-context prompt of the form $a_{i_{1}}b_{i_{1}}a_{i_{2}}b_{i_{2}}\cdots a_{i_{\ell}}$ with distinct $i_{j}$’s, the model never predicts $b_{i_{\ell}}$: it predicts a uniform probability vector over $\mathcal{V}$ when $1\leq\ell\leq k$, and $b_{i_{1}}$ when $\ell=k+1$.
Results. Theorem 7 highlights the finding that the success of ICL relies heavily on how the patterns appear in the training data. In this scenario, the $(a_{i},b_{i})$ pairs consistently appear at the beginning and end of each training sentence, and we anticipate the model to recognize this relationship for ICL to occur. However, as shown in Theorem 7 and Table 7, this is not the case.
@@ 4.3 Experiment on a synthetic corpus
We conduct an experiment on a synthetic corpus featuring (country)-(capital) relationships. Each sentence falls into one of four categories: (1) exactly one country-capital pair, (2) exactly two country-capital pairs, (3) a single country without a pair, and (4) no country. In sentences with one country-capital pair, the capital appears in the first position, the country in the last, and each sentence contains six words (as in Section 4.2). The corpus generation process is detailed in Appendix J.
@@
We train a five-layer two-head autoregressive transformer on this corpus, with an embedding dimension of $100$. Similar to Section 2.4, we evaluate the ICL accuracies using prompts involving countries and their capitals. The results show zero ICL accuracy across varying in-context examples (one to five), supporting our theory.
Discussion
This paper examines how in-context learning (ICL) arises from pre-training on unstructured language data, with three key findings: (1) ICL for word analogy tasks can emerge from simple co-occurrence modeling, using models like continuous bag of words (CBOW) without positional encoding or attention; (2) positional information and structured nuisance tokens are essential for ICL in logic reasoning tasks that require recognizing rare patterns and generalizing to new tokens; and (3) the structure of training data significantly impacts ICL effectiveness.
Limitations and future work
This study has several limitations. Firstly, the experiments are conducted on a relatively small scale. However, they still provide sufficient evidence to support the theoretical findings. Secondly, the focus of this study is on specific types of in-context learning (ICL) tasks, as described in Section 1. Thirdly, the pre-training data considered in this work may not match the valid grammatical sentences that language models are usually trained on. Nonetheless, our co-occurrence results still apply to grammatical sentences, as the co-occurring pairs can appear naturally within them (e.g., "Beijing is the capital of China," or "the city of Beijing is located in China"). Lastly, real data sets are not utilized due to the lack of alignment with the study objectives.
Despite these limitations, we believe this work provides valuable understanding of the key factors enabling ICL to occur from training on unstructured natural language data, supported by both theoretical and empirical evidence from experiments involving prompting and synthetic data. Further analyses on other ICL tasks and their reliance on model architecture can be fruitful avenues for future work.
Acknowledgements. This work was supported in part by the Office of Naval Research under grant number N00014-23-1-2590, the National Science Foundation under grant numbers 2231174 and 2310831, No. 2428059, and a Michigan Institute for Data Science Propelling Original Data Science (PODS) grant.
Supplementary Material
Appendix A Related work
@@
Large language models (LLMs), such as transformers, are widely recognized for their outstanding performance in in-context learning (ICL) [10]. ICL refers to the capability of LLMs to discern specific tasks and generate predictions based on prompt exemplars without needing any parameter updates. A multitude of studies have been dedicated to exploring this intriguing phenomenon from various theoretical and empirical perspectives. In this section, we provide a brief summary of some of these studies.
@@
Some studies adopted a Bayesian approach to studying ICL. Xie et al. [60] posited that ICL can be viewed as implicit Bayesian inference. They demonstrated that LLMs can infer a latent document-level concept for next-token prediction during pre-training and a shared latent concept across input-output pairs in an ICL prompt, under the assumption that documents are generated from hidden Markov models (HMMs). Wang et al. [56] and Zhang et al. [67] expanded on this idea by exploring more realistic latent variable models beyond HMMs. Wang et al. [56] argued that large language models function as latent variable models, with latent variables containing task-related information being implicitly inferred. Zhang et al. [67] showed that without updating the neural network parameters, ICL can be interpreted as Bayesian model averaging parameterized by the attention mechanism. Panwar et al. [39] provided empirical evidence that transformers behave like Bayesian predictors when performing ICL with linear and non-linear function classes. Dalal and Misra [18] proposed a Bayesian learning framework to understand ICL through the lens of text generation models represented by multinomial transition probability matrices. Chiang and Yogatama [14] proposed the pelican soup framework to explain ICL without relying on latent variable models. This framework incorporates concepts such as a common sense knowledge base, natural language classification, and meaning association, enabling the establishment of a loss bound for ICL that depends on the number of in-context examples.
@@
In a related exploration, Li et al. [31] showed that softmax regression models learned through gradient descent are similar to transformers. Ren and Liu [44] related ICL with softmax transformers to contrastive learning, where the inference process of ICL can be viewed as a form of gradient descent. Mahankali et al. [35] proved that minimizing the pre-training loss is equivalent to a step of gradient descent in single-layer linear transformers. Vladymyrov et al. [54] established that linear transformers execute a variant of preconditioned gradient descent by maintaining implicit linear models. On the other hand, some studies argued that the ICL ability of transformers cannot be attributed to gradient descent. Fu et al. [20] showed that ICL for linear regression tasks arises from higher-order optimization techniques like iterative Newton’s method rather than gradient descent. Wibisono and Wang [57] demonstrated that transformers can perform ICL on unstructured data whose prompt exemplars lack explicit pairings, with softmax attention playing an important role especially when using a single attention layer. Shen et al. [46] provided empirical evidence that the equivalence between gradient descent and ICL might not be applicable in real-world scenarios. In contrast to these studies, our work provides a connection between ICL and classical language models like continuous bag of words (CBOW). Specifically, we show that ICL for word analogy tasks with semantically related word pairs can arise by modeling co-occurrence patterns via CBOW.
@@
Numerous studies focused on the pre-training aspects (e.g., data distribution and task diversity) of ICL. Min et al. [37] showed that the input-label mapping in the in-context examples does not significantly affect ICL performance. Chan et al. [11] demonstrated that the ICL capabilities of transformers depend on the training data distributions and model features. Kossen et al. [30] established that ICL considers in-context label information and is capable of learning entirely new tasks in-context. Li and Qiu [32] introduced an iterative algorithm designed to enhance ICL performance by selecting a small set of informative examples that effectively characterize the ICL task. Qin et al. [41] proposed a method based on zero-shot chain-of-thought reasoning for selecting ICL examples, emphasizing the importance of choosing diverse examples that are strongly correlated with the test sample. Han et al. [25] studied ICL by identifying a small subset of the pre-training data that support ICL via gradient-based methods. They discovered that this supportive pre-training data typically consist of more uncommon tokens and challenging examples, characterized by a small information gain from long-range context. Peng et al. [40] proposed a selection method for ICL demonstrations that are both data-dependent and model-dependent. Van et al. [52] introduced a demonstration selection method that enhances ICL performance by analyzing the influences of training samples using influence functions.
@@
In a similar vein, Wu et al. [59] demonstrated that pre-training single-layer linear attention models for ICL on linear regression with a Gaussian prior can be effectively accomplished with a minimal number of independent tasks, regardless of task dimension. Raventós et al. [42] emphasized a task diversity threshold that differentiates the conditions under which transformers can successfully address unseen tasks. Yadlowsky et al. [63] attributed the impressive ICL capabilities of transformers to the diversity and range of data mixtures in their pre-training, rather than their inductive biases for generalizing to new tasks. Ding et al. [19] compared the ICL performance of transformers trained with prefixLM (where in-context samples can attend to all tokens) versus causalLM (where in-context samples cannot attend to subsequent tokens), finding that the latter resulted in poorer ICL performance. Chen et al. [13] discovered that the ICL capabilities of language models rely on the presence of pairs of phrases with similar structures within the same sentence. Zhao et al. [68] proposed a calibration scheme that modifies model parameters by adding random noises, resulting in fairer and more confident predictions. Abbas et al. [1] demonstrated that the ICL predictions from transformer-based models often exhibit low confidence, as indicated by high Shannon entropy. To address this issue, they introduced a straightforward method that linearly calibrates output probabilities, independent of the model’s weights or architecture. Similar to these works, our work highlights the importance of co-occurrence, positional information, and training data structure for ICL to arise.
@@
A number of works delved into specific data generating processes to provide insight into the emergence of ICL. Bhattamishra et al. [8] examined the ICL ability of transformers by focusing on discrete functions. Specifically, they showed that transformers perform well on simpler tasks, struggle with more complex tasks, and can learn more efficiently when provided with examples that uniquely identify a task. Guo et al. [22] investigated ICL in scenarios where each label is influenced by the input through a potentially complex yet constant representation function, coupled with a unique linear function for each instance. Akyürek et al. [6] studied ICL of regular languages produced by random finite automata. They compared numerous neural sequence models and demonstrated that transformers significantly outperform RNN-based models because of their ability to develop n-gram heads, which are a generalization of induction heads. Sander et al. [45] analyzed simple first-order autoregressive processes to gain insight into how transformers perform ICL to predict the next tokens. On the other hand, our work focuses on data generating processes containing semantically related word pairs and repeated token patterns to better understand several components that are crucial for ICL to occur from training on unstructured data.
@@
Some studies explored how different components of transformers affect their ICL abilities. Ahuja and Lopez-Paz [4] compared the ICL performance of transformers and MLP-based architectures under distribution shifts. Their findings demonstrate that while both methods perform well in in-distribution ICL, transformers exhibit superior ICL performance when faced with mild distribution shifts. Collins et al. [15] showed that softmax attention outperforms linear attention in ICL due to its ability to calibrate its attention window to the Lipschitzness of the pre-training tasks. Xing et al. [61] focused on linear regression tasks to identify transformer components that enable ICL. They found that positional encoding is crucial, along with the use of multiple heads, multiple layers, and larger input dimensions. Cui et al. [16] proved that multi-head attention outperforms single-head attention in various practical scenarios, including those with noisy labels and correlated features. Chen et al. [12] investigated the ICL dynamics of a multi-head softmax attention model applied to multi-task linear regression. They proved the convergence of the gradient flow and observed the emergence of a task allocation phenomenon, where each attention head specializes in a specific task.
@@
Finally, several studies proposed various hypotheses on the emergence of ICL and provided theoretical justifications. Swaminathan et al. [49] introduced clone-structured causal graphs (CSCGs) to explain how ICL can generalize to unseen sentences via a mechanism called rebinding. Li et al. [33] viewed ICL as an algorithm learning problem where a transformer implicitly constructs a hypothesis function at inference time. Han et al. [24] argued that the ability of transformers to execute ICL is attributable to their capacity to simulate kernel regression. Singh et al. [47] explored the interaction between ICL and in-weights learning (IWL) using synthetic data designed to support both processes. They observed that ICL initially emerges, followed by a transient phase where it disappears and gives rise to IWL. Yan et al. [64] studied ICL from the perspective that token co-occurrences play a crucial role in guiding the learning of surface patterns that facilitates ICL. Abernethy et al. [2] showed that transformers can execute ICL by dividing a prompt into examples and labels, then employing sparse linear regression to deduce input-output relationships and generate predictions. Lin and Lee [34] developed a probabilistic model that can simultaneously explain both task learning and task retrieval aspects of ICL. Here, task learning refers to the ability of language models to identify a task from in-context examples, while task retrieval pertains to their ability to locate the relevant task within the pre-training data.
Appendix B Proof of Theorem 1